OpenStudy (anonymous):

Find the volume of a rectangular prism that has a length of 10 inches, width of 5 inches, and height of 6 inches

OpenStudy (anonymous):

Volume = length * width * height = 10 in * 5 in * 6 in 5*6 = 30 30*10=300 then don't forget to keep track of your units: in*in*in = in^3 so 300 in^3

OpenStudy (anonymous):

Thanks :)